OpenStudy (anonymous):

How many GRAMS of copper(2) iodide are present in 4.95 moles of this compound? How many MOLES of copper(2) iodide are present in 4.73 grams of this compound?

OpenStudy (anonymous):

Would this not be better asked in Chemistry? They're used to dealing with this sort of stuff... Anyway \[n \times RMM = m\] RMM of Cu(II)I2 is I think: 63.15+126.9+126.9= 316.95 Although copper 2 iodide isn't stable - but anyway! 4.95 x 316.95 = 1,568.9025g (or 1.57kg) Same again but backwards: n = m/RMM n = 4.73/316.95 n = 0.01492